OpenStudy (anonymous):

5(3x+2)-4x=17-(6-4x) show work

OpenStudy (anonymous):

Start by distributing the 5 into the (3x+2) to get 15x+10. Now you have 15x+10-4x=17-(6-4x). Next step would be to distribute the negative 1 into the (6-4x) to get -6+4x. You will then have. 15x+10-4x=17-6+4x. Now all you have to do is combine like terms on both sides. Giving you 11x+10=11+4x. Next step would be getting all of you x terms to the left side. giving you 7x+10=11. After doing this you subtract 10 from 11 to get. 7x=1 and for the last step you just divide both sides by 7 to get x=1/7

OpenStudy (anonymous):

thank !!!